if x~bin 7,1/4 find p(x=6)

if x~bin 7,1/4 find p(x=6)
Answer #1

Answer:

Given,

n = 7

p = 1/4

= 0.25

q = 1 - 0.25

= 0.75

Here let us consider,

Binomial distribution

P(X = r) = nCr*p^r*q^(n-r)

P(X = 6) = 7C6*0.25^6*0.75^(7-6)

= 7*0.25^6*0.75

= 0.0013

P(x = 6) = 0.0013